A manufacturing operation produces products that contain unacceptable coloring 30% of the time. What is the probability that a sample of 3 products will contain 0 that have unacceptable coloring?
Answer #1

P(unacceptable coloring) = 0.30

P(acceptable coloring) = 1 - 0.30 = 0.70

P(3 products will contain 0 that have unacceptable coloring) = P(all 3 have acceptable coloring)

= 0.73

= 0.343